OpenStudy (anonymous):

what is the energy in joules of one quantum of 55.4 nm reply in scientific notation

OpenStudy (xishem):

\[c= \lambda \nu \rightarrow \nu = \frac{c}{\lambda}\]Where... c is the speed of light lambda is the wavelength nu is the frequency\[\nu = \frac{3.00*10^8m*s^{-1}}{55.4nm}*\frac{10^9nm}{1m}=5.4152*10^{15}Hz\] \[E=h \nu\]Where... E is the energy of the quanta h is Planck's constant nu is the frequency of the quanta \[E=(6.626*10^{-34}\frac{m^2*kg}{s})(5.4152*10^{15}\frac{1}{s})\]Simplifying all the junk on the right, we end up getting...\[E=3.59*10^{-18}\frac{m^2*kg}{s^2}=3.59*10^{-18}J\]